OpenStudy (anonymous):

the question is What is the circumference of the circle? the radius is 8 inches (or diameter is 16 inches). choices are a. 8 pi inches; b. 16 pi inches; c. 24 pi inches; or d. 48 pi inches. I say it's 16 pi inches and my boys (incoming 7th grade advanced math students) are saying d. 48 pi inches

OpenStudy (anonymous):

Circumference = d * pi Answer = 16pi inches

OpenStudy (wolf1728):

Circumference = diameter * PI Circumference = 16 * PI

OpenStudy (anonymous):

They still don't believe me so I've emailed their math teacher from last year. Thanks.

OpenStudy (anonymous):

Here is the actual question 2.

OpenStudy (anonymous):

ok... It's B